a)What is the mass of heavy water D2O produced when 5.20g of O2 reacts with excess D2?
b)What is the mass of water H2O,produced when 5.20g of O2 reacts with excess H2

Consider the reaction 2D2+O2->2D2O

Can someone step me through this?
Answer #1
Look at it from moles only.
how many moles of O2 is in 5.2grams?
mole=5.2/32 work this precisely.
YOU will get twice this moles of product. What is the mass?
Moles product: 5.2/32 * 2

so what is the mass of the product:
massD2O= 5/32 *2*molmassD2O
massH2O=5/32 * 2 * molmassH2O
answered by: Shakaa'